PROBLEM D: On January 1, 2020 Lance Co. issued five-year bonds with a face value of $1,000,000 and a stated interest rate ofProblem D
Answer #1
a PV of Interest annuity $60,000*PVAF(5%,10 years) =$60,000*7.72173 =$463,304
(Interest on Bond =$1,000,000*12%*6/12 =$60,000)
b PV of principal $1,000,000*PVIF(5%,10 years) =$1,000,000*0.61391 =$613,910
c Issue Price($463,304 + $613,910) $1,077,214
d Premium on Bond issued =$1,077,214 - $1,000,000 =$77,124
Accounts and explanation Debit(in $) Credit(in $)
Cash                                         10,77,214
     Bonds Payable                    10,00,000
    Premium on Bonds Payable                         77,214
(To Bond issued at premium)